A person deposited $1000 at the end of each year into a retirement savings plan for the last 10 years. Her deposits earned interest at 8% for the first 3 years; at 10.25% for the next 4 years; and at 9% for the last 3 years. (a) what is the accumulated value of her retirement plan? (b) What is the total interest earned for the 10 years?
Answer by ptfile(81)

Let P = Principal
    R = Interest rate per annum
    T = Duration of the loan
    I = Interest
   FV = Future Value
   PV = Present Value

a) 
First three years
FV = PV (1 + r)^t
FV = 1000 (1 + .08)^3
FV = 1259.71

Next four years
FV = PV (1 + r)^t
FV = 1259.71 (1 + .1025)^4
FV = 1861.17

Last three years
FV = PV (1 + r)^t
FV = 1861.17 (1 + .09)^3
FV = 2410.27

b)
2410.27 - 1000 = 1410.27

The amount she accumulated in 10 years is $2410.27 and the total interest she earned in 10 years is $1410.27.